22and 1fourth multiply 4and 1third =
Answer by cathieb1225(41)

22and 1fourth multiply 4and 1third =
(22 1/4)(4 1/3)
Convert both into improper fractions. You do this by multiplying the number with the denominator and adding the numerator. Thus, 22 1/4=89/4 and 4 1/3=13/3
(89/4)(13/3)=1157/12=96 5/12